OpenStudy (batmaaaanxd):

Find the condition that the system of equations ax+by= c, lx+my= n has a unique solution?

OpenStudy (ageta):

http://prntscr.com/apokps

OpenStudy (phi):

do you know about matrices?

OpenStudy (phi):

\[\left[\begin{matrix}a & b \\ l & m\end{matrix}\right] \left[\begin{matrix}x \\ y\end{matrix}\right]=\left[\begin{matrix}c \\ n\end{matrix}\right] \]

OpenStudy (phi):

to solve you multiply both sides by the inverse of the matrix you get \[ \left[\begin{matrix}x \\ y\end{matrix}\right]= \frac{1}{am-bl}\left[\begin{matrix}m & -b \\ -l & a\end{matrix}\right]\left[\begin{matrix}c \\ n\end{matrix}\right] \] and that has a solution as long as (am-bl) does not equal 0 in other words, the condition is \[ am \ne bl \]

OpenStudy (batmaaaanxd):

@phi Nope, I do not know

OpenStudy (phi):

remember in your previous problem where you wrote the coefficients of x,y <a,b> and <l,m> in this case if there is a multiplier, (call it "g") so that ga = l and gb= m (in your problem we found ½ worked) then that makes the two lines parallel if we have ga = l then g= l/a and similarly gb= m means g= m/b so if g exists, it requires l/a = m/b or cross-multiplying, bl = am and fog g to not exist, we want \(bl \ne am\)

OpenStudy (phi):

*and for g not to exist, we want \( bl \ne am\)